40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Jun 7 10:26:36 2018, keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. photodetector_alignment.jpgdesign1.PNG
    Reply  Thu Jun 7 22:36:09 2018, not keerthana, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 00:36:54 2018, gautam, Update, PSL, observing the resonance signal corresponding to the injected frequency. 
    Reply  Fri Jun 8 01:08:22 2018, johannes, Update, PSL, First light of AUX at YEND AG4395A_07-06-2018_205019.pdfPSL_AUX_SETUP.pdfAS_AUX_SETUP.pdf
       Reply  Mon Jun 18 14:32:42 2018, gautam, Update, PSL, Optics on AS table 
          Reply  Mon Jun 18 15:59:17 2018, johannes, Update, PSL, Optics on AS table 
             Reply  Mon Jun 18 16:57:54 2018, Koji, Update, PSL, Optics on AS table 
             Reply  Fri Jun 22 14:18:04 2018, Steve, Update, PSL, Optics on AS table 
Message ID: 13932     Entry time: Fri Jun 8 01:08:22 2018     In reply to: 13924     Reply to this: 13981
Author: johannes 
Type: Update 
Category: PSL 
Subject: First light of AUX at YEND 

Among the things that we hadn't taken care of yesterday before beginning to look for transmission signals were the polarization of the AUX beam on the AS table and optimizing the PLL feedback. The AUX beam is s-polarized on the PSL table (choice due to availablility of mirrors), and I added a half waveplate in front of the fiber to match it's axes. I placed another half-waveplate at the fiber output and send the reflection port of a PBS cube onto a PDA1CS photodetector. By alternatingly turning the waveplates I minimized the reflected light, giving strongly p-polarized light on the AS table for best results when interfering with the IFO beam. I wiggled the fiber and found no strong dependency of the output polarization on fiber bending. Attachment 2 shows the current layout.

The beat signal between AUX and PSL table is at -20dBm, and I adjusted the PLL gain and PI-corner to get reliable locking behavior. I think it's a good idea to keep the AUX beam on the AS table blocked while it's not in use, and only unblock it when it is phaselocked to avoid a rogue beam with no fixed phase relation to the PSL in the IFO.I blocked the beam after completing this work today.

I used the signal chain that Keerthana, Koji, and I set up yesterday to look for mode flashed of the AUX light in the YARM using the RF beat with the PSL carrier in transmission. To align the AUX beam to the arm the following steps were performed:

  1. Using a spectrum analyzer to look at the RF power at the target frequency between frequency-shifted AUX beam and PSL carrier on AS110, align the beam using the mirror pair closest to the fiber coupler for maximum signal.
  2. Initiate a sweep of the PLL LO frequency sourced by the Marconi using GPIB scripts over about 1 FSR. A strong peak was visible at ~31.76 MHz offset frequency
  3. Tune and hold LO frequency (in this case at 48.2526 MHz) such that AUX beam resonates in the arm. Optimize alignment by maximizing RF signal on PD in transmission.

This was followed by a sweep over two full FSRs. Attachment #1 shows the trace recorded by the AG4395 using the max data hold setting during the sweep. Essentially the beat between AUX and PSL carrier traced out the arm's transmission curve. At minimum transmission there was still a ~82dB beat on the transmission PD visible.

The YEND QPD is currently blocked and sees no light.

Attachment 1: AG4395A_07-06-2018_205019.pdf  22 kB  Uploaded Fri Jun 8 02:44:35 2018  | Hide | Hide all
AG4395A_07-06-2018_205019.pdf
Attachment 2: PSL_AUX_SETUP.pdf  4.086 MB  Uploaded Fri Jun 8 02:51:50 2018  | Hide | Hide all
PSL_AUX_SETUP.pdf
Attachment 3: AS_AUX_SETUP.pdf  2.065 MB  Uploaded Fri Jun 8 02:52:16 2018  | Hide | Hide all
AS_AUX_SETUP.pdf
ELOG V3.1.3-